# Selection of Filters for Photonic Crystal Spectrometer Using Domain-Aware Evolutionary Algorithms

### v1

@@ 4.1.4 Noisy Portfolio
This algorithm corresponds to the method called RecombiningPortfolioOptimisticNoisyDiscreteOnePlusOne in the library [44]. According to the documentation in the source, this algorithm constitutes an extension of $(1+1)$ EA. The extension includes genetic crossover applied at every iteration and a so-called optimistic noise handling scheme. One can find more details about those methods in the source code of Nevergrad [44]. Moreover, the mutation rate in this algorithm is controlled during the optimization process using uniform mixing proposed in [47].
@@ 4.2 Response Surface Methodologies
Response Surface Methodology (RSM) is an optimization algorithm, where a model, called response surface, approximates the behavior of the objective function based on a set of known input-output data points [23]. RSM starts optimization by creating a Design of Experiment (DoE), meaning a set of uniformly selected points that are used to build the initial model. Then RSM iteratively uses the surface as a surrogate to find the next infill point, meaning the point where the objective value is evaluated. The infill point is then added to the model. Those iterations are repeated until the computational budget is exhausted.
4.2.1 BO
Bayesian Optimization (BO) [48] and Efficient Global Optimization (EGO) [49] is a particular case of RSM, where the response surface constitutes a statistical model. We considered the conventional Gaussian Process (GP) [50] as the probabilistic model. This model is especially appealing for our application because it allows modeling the inherent Gaussian noise in the objective function. The time complexity of the algorithm that creates GP is $O(n^{3})$, where $n$ is the number of points added to the model [51]. It becomes computationally inefficient to add more than $10^{3}$ points to GP, therefore we limit the computational budget for BO by $10^{3}$ evaluations of $\widehat{\mathrm{F}}_{K}$. We used BO implementation from skopt library [52]. The number of the DoE points was set to $110$, and the noise parameter of the algorithm was set to Gaussian, due to the reasons discussed in Section 3.2.2. For each of the $M$ components, we defined the bounding box, which restricts the search space for BO, to be $[1,L]$. All float values that are encountered in the infill points are converted to integers by disregarding the decimal parts. The rest of the parameters of BO were set to their default values used in skopt.
@@ 4.3 Model-Based Methods
Model-based methods maintain probabilistic models to navigate in the search space. These models are built and refined iteratively to guide the search towards promising regions. Here we consider two algorithms that fall under this category. Generally speaking, those algorithms could also be classified as Evolutionary Algorithms. Notably, those optimizers are different from the ones described in 4.1 because the model-based methods learn and adapt their inherent models throughout the optimization process.
4.3.1 $(\mu/2,\lambda)$ MIES
Mixed Integer Evolution Strategy [53] is proposed to address the problems with domains that contain real and discrete parts. The probabilistic model is used in this algorithm to sample mutations. Since in our OFS problem filters are encoded as integers, we apply only part of the algorithm that addresses integer domains. For such search spaces, MIES borrows the method originally proposed for unbounded sets of values in every component [54]. We used the implementation of this algorithm from the library [55]. The class that corresponds to MIES is called mipego.optimizer.mies.MIES. The selection scheme used in the algorithm is non-elitist, meaning only the created $\lambda$ individuals take part in the selection. Before the mutation, MIES applies uniform crossover to a pair of parents selection uniformly at random from the population of $\mu$ solutions. In our experiments, we used values $\mu=10,\lambda=20$.
4.3.2 UMDA
1 For every $i\in\overline{1,M},j\in\overline{1,L}$ initialize: $\mathrm{P}^{(1)}_{i,j}\leftarrow 1/L$;
@@ 18             Update best-so-far: $\mathbf{x}^{*}\leftarrow\mathbf{y}^{(1)}$;
Algorithm 1 UMDA for integer programming, where objective function $F:\mathbb{L}^{M}\to\mathbb{R}$ is minimized. Points in its domain have dimensionality $M$. Value at every component of every point is an integer from the set $\mathbb{L}\coloneqq\overline{1,L}$. Given positive integers $\mu\leq\lambda$, the algorithm produces $\lambda$ offspring and selects $\mu$ ones from them. The maximal number of $F$ evaluations is limited by constant $b$.
Estimation of Distribution Algorithms (EDAs) [56, 57] explicitly define the probabilistic distribution over the search space of an optimization problem. Univariate Marginal Distribution Algorithm (UMDA) [58] is an instance of EDA suited for problems with pseudo-boolean search spaces. UMDA maintains a global probability distribution over the search space, considering the value in every component as a random variable. The core assumption behind UMDA is that those random variables are independent. UMDA is an iterative optimization algorithm. On every iteration, it samples $\lambda$ candidate solutions from the global distribution and uses the first $\mu$ solutions in the ranking according to the objective function to update this probabilistic model. Since usually the algorithm is considered only for pseudo-boolean domains, we propose our version of UMDA for integer domains in Algorithm 1. In our implementation, we used constants $\mu=5,\lambda=50$.
@@ 4.4 Automated Algorithm Selection
The optimizer may be suggested for a given problem using Automated Algorithm Selection (AAS) [59]. Usually, AAS methods are trained for the specific instances of optimization problems and then applied to previously unseen tasks.
4.4.1 NGOpt
NGOpt is a so-called wizard, which selects an optimization algorithm based on the properties of the considered problem and can switch the algorithm during the minimization process [60]. The properties include the number of dimensions, presence of noise, type of constraints, etc. We used NGOpt implemented in the Nevergrad platform.
@@ 4.5 Numerical Comparison
In this section, we demonstrate the application of the following selected algorithms to the OFS problem. The selected algorithms are: NGOpt (Section 4.4.1), $(\mu/2+\lambda)$ MIES (Section 4.3.1), Integer Fast GA (Section 4.1.3), Noisy Portfolio (Section 4.1.4), BO (Section 4.2.1), $(\mu+\lambda)$ EA (Section 4.1.1), $(\mu/2+\lambda)$ EA (Section 4.1.2) and UMDA (Section 4.3.2). Let us denote the set of those algorithms as $\mathbb{A}_{1}$. We used $\widehat{\mathrm{F}}_{K}$ as the objective function with fixed constant $K=10^{3}$ for all the algorithms. In each algorithm except BO, we used the computational budget of $2.1\cdot 10^{3}$ evaluations of the objective function $\widehat{\mathrm{F}}_{K}$. For BO we made only $10^{3}$ evaluations of $\widehat{\mathrm{F}}_{K}$, due to the reasons discussed in Section 4.2.1. Every algorithm is run $n=20$ times and every time the set of initial solutions is taken uniformly at random from the search space $\mathscr{L}$.
@@
    Quality of the solution considered by algorithm $A$ when computational budget $t$ is spent:
@@
    Quality of best-so-far solution found by algorithm $A$ when computational budget $t$ is spent:
@@
The results of this experiment are summarized in Figure 7. In both plots on this figure, the X axis depicts the number of times $t$ function $\widehat{\mathrm{F}}_{K}$ was evaluated. The Y axis in part (a) of Figure 7 represents the approximated mean of $\mathrm{f}$ criterion. This plot shows that candidate solutions in Integer Fast Ga, Noisy Portfolio, and BO vary and do not improve on average when the optimization process proceeds. However, NGOpt starts producing solutions that perform better on average when approximately half of the given computational budget is exhausted. This can happen due to the switch of the underlying optimization algorithm that NGOpt uses. Average candidate solutions considered by $(\mu/2,\lambda)$ MIES significantly improve in the first half of the optimization process, but then stagnate until the end.
The Y axis in part (b) of Figure 7 represents the approximated mean of $\mathrm{g}$ criterion and the shaded area represents the approximated standard deviation of $\mathrm{g}$ for each algorithm in $\mathbb{A}_{1}$. The line for every algorithm in this plot constitutes the lower envelope of the corresponding line in the plot (b) of Figure 7. All the algorithms except $(\mu/2,\lambda)$ MIES on average produced some solutions that outperform the baseline solution. It is clear, that algorithm $(\mu+\lambda)$ EA, $(\mu/2+\lambda)$ EA and UMDA outperform all the other considered algorithms in both criteria $\mathrm{f}$ and $\mathrm{g}$. We call those three algorithms leading algorithms and aim to improve their performance in our further investigation.
@@ 5 Employing New Distances to Navigate in the Domain of OFS
When an optimization algorithm is equipped with domain information it can efficiently disregard areas of the search space with sub-optimal candidate solutions. While there are many possible ways to account for problem-specific information, we use a distance metric as proposed in [61]. The metric represents our knowledge of the filters in the library $\mathbb{L}$ and the search space symmetry to permutations.
5.1 Metric between candidate solutions
In this section, we propose two distance metrics on the space $\mathscr{L}$. We remind the following classical definition of the metric space in Definition 1.
Definition 1 (Metric Space).
  1. 1.
    $d(x,y)\geq 0$ and $d(x,y)=0\iff x=y$,
  2. 2.
    $d(x,y)=d(y,x)$,
  3. 3.
    $d(x,z)\leq d(x,y)+d(y,z)$.
The function $d$ is called the metric on $X$.
@@ 5.1 Metric between candidate solutions
$d_{2}(x,y)=\left|\sum_{\zeta}\zeta^{2}\tilde{T^{x}_{\xi}}-\sum_{\zeta}\zeta^{2}\tilde{T^{y}_{\zeta}}\right|.$ (17)
A larger second moment represents a filter with sharper and or more spectral features. We numerically validated that there are no different filters $x,y\in\mathbb{L}$ such that $d_{1}(x,y)=0$ or $d_{2}(x,y)=0$. The other conditions of the metric are satisfied for $d_{1},d_{2}$ due to the properties of the modulo. Therefore, $d_{1},d_{2}$ are metrics on the space $\mathbb{L}$.
@@
Any two points $\bm{x},\bm{y}\in\mathbb{L}^{M}$ are in relation $\bm{x}\equiv\bm{y}$ if $\textsc{LAP}(d_{\mathbb{L}},\bm{x},\bm{y})=0$. This relation induces the quatient space $\mathbb{M}(d_{\mathbb{L}})\coloneqq\mathbb{L}^{M}/\equiv.$
Proposition 1.
Proof.
  1. 1.
  2. 2.
  3. 3.
    There exist bijections $\pi_{1},\pi_{2}\in\mathbb{P}$ such that:
    $\mathscr{D}(\bm{x},\bm{y})+\mathscr{D}(\bm{y},\bm{z})$
    $\geq\sum_{i=1}^{k}d_{\mathbb{L}}(x_{i},z_{\pi_{2}\circ\pi_{1}(i)})$
@@ 5.2 Distance-Driven Versions of $(\mu+\lambda)$ EA and $(\mu/2+\lambda)$ EA
We apply the method proposed by us earlier in [61] to employ the distance metric $d$ in EA. It extends a mutation operator in the optimization algorithm, which aims to find at least one $\bm{x}^{*}\in\mathbb{X}$ such that $F(\bm{x}^{*})=\min\nolimits_{\bm{x}\in\mathbb{X}}\left\{F(\bm{x})\right\},$ where $\mathbb{X}$ is some search space and $F$ is the objective function defined on it. In order to make this extension, the method defines a mapping of the distance between any two candidate solutions to the interval $(0,1)$ and creates a distribution $\mathcal{D}$ over the step-sizes on this interval. To do a mutation of the solution $\bm{x}\in\mathbb{X}$, the algorithm samples a step size $s\in(0,1)$ from the distribution $\mathcal{D}$ and maps this step size back to the distance $S\in\mathbb{R}$ in the original domain $\mathbb{X}$. Then it solves the internal optimization problem to generate a solution $\bm{y}\in\mathbb{X}$ which steps away from $\bm{x}$ to the distance $S$:
@@
We explain how we applied every step (emphasized with italic font) of this general methodology in later sections: Section 5.2.1, Section 5.2.2, and Section 5.2.3. Now we directly describe the proposed extension of $(\mu/2+\lambda)$ EA summarized in Algorithm 2. The extension of $(\mu+\lambda)$ EA is the same, except that instead of the lines 2, 2, and 2 it chooses $\mathbf{c}$ uniformly at random from $\mathfrak{P}^{(t)}.$
1 Select points for exploration for each $i\in\overline{1,R}$: $\mathbf{e}^{(i)}\leftarrow\textsc{u.a.r.}\!\left(\mathbb{L}^{M}\right)$;

### v2

@@ 4.1.4 Noisy Portfolio
This algorithm corresponds to the method called RecombiningPortfolioOptimisticNoisyDiscreteOnePlusOne in the library Nevergrad [40]. According to the documentation in the source, this algorithm constitutes an extension of $(1+1)$ EA. The extension includes genetic crossover applied at every iteration and a so-called optimistic noise handling scheme. One can find more details about those methods in the source code of Nevergrad [40]. Moreover, the mutation rate in this algorithm is controlled during the optimization process using uniform mixing proposed in [43].
@@ 4.2 Response Surface Methodologies
Response Surface Methodology (RSM) is an optimization algorithm, where a model, called response surface, approximates the behavior of the objective function based on a set of known input-output data points [19]. RSM starts optimization by creating a Design of Experiment (DoE), meaning a set of points (for example uniformly selected), which is used to build the initial model. Then RSM iteratively uses the surface as a surrogate to find the next point, where the objective value is evaluated. This point is then added to the model. Those iterations are repeated until the computational budget is exhausted.
4.2.1 Bayesian Optimization
Bayesian Optimization (BO) [44] and Efficient Global Optimization (EGO) [45] are particular cases of RSM, where the response surface constitutes a statistical model. We considered the conventional Gaussian Process (GP) [46] as the probabilistic model. This model is especially appealing for our application because it allows modeling the inherent Gaussian noise in the objective function. The time complexity of the algorithm that creates GP is $O(n^{3})$, where $n$ is the number of points added to the model [47]. It becomes computationally inefficient to add more than $10^{3}$ points to GP, therefore we limit the computational budget for BO by $10^{3}$ evaluations of $\widehat{\mathrm{F}}_{K}$. We used BO implementation from skopt library [48]. The number of the DoE points was set to $110$, and the noise parameter of the algorithm was set to Gaussian, due to the reasons discussed in Section 3.2.1. For each of the $M$ components, we defined the bounding box, which restricts the search space for BO, to be $[1,L]$. All float values that are encountered in the evaluated points are converted to integers by disregarding the decimal parts. The rest of the parameters of BO were set to their default values used in skopt.
@@ 4.3 Model-Based Methods
Model-based methods maintain probabilistic models to navigate in the search space. These models are built and refined iteratively to guide the search towards promising regions. Here we consider two algorithms that fall under this category. Generally speaking, those algorithms could also be classified as Evolutionary Algorithms. Notably, those optimizers are different from the ones described in Section 4.1 because the model-based methods learn and adapt their inherent models throughout the optimization process.
4.3.1 $(\mu/2,\lambda)$ Mixed Integer Evolution Strategy
Mixed Integer Evolution Strategy (MIES) [49] is proposed to address the problems with domains that contain real and discrete parts. The probabilistic model is used in this algorithm to sample mutations. Since in our OFS problem filters are encoded as integers, we apply only part of the algorithm that addresses integer domains. For such search spaces, MIES borrows the method originally proposed for unbounded sets of values in every component [50]. We used the implementation of this algorithm from the library [51]. The selection scheme used in the algorithm is non-elitist, meaning only the created $\lambda$ individuals take part in the selection. Before the mutation, MIES applies uniform crossover to a pair of parents selection uniformly at random from the population of $\mu$ solutions. In our experiments, we used values $\mu=10,\lambda=20$.
4.3.2 Univariate Marginal Distribution Algorithm
1For every $i\in\overline{1,M},j\in\overline{1,L}$ initialize: $\mathrm{P}^{(1)}_{i,j}\leftarrow 1/L$;
@@ 18      Update best-so-far: $\mathbf{x}^{*}\leftarrow\mathbf{y}^{(1)}$;
Algorithm 1 UMDA for integer programming, where objective function $F:\mathbb{L}^{M}\to\mathbb{R}$ is minimized. Value at every component of every point is an integer from the set $\mathbb{L}\coloneqq\overline{1,L}$. Given positive integers $\mu\leq\lambda$, the algorithm produces $\lambda$ offspring and selects $\mu$ ones from them. The maximal number of $F$ evaluations is limited by constant $b$.
Estimation of Distribution Algorithms (EDAs) [52, 53] explicitly define the probabilistic distribution over the search space of an optimization problem. Univariate Marginal Distribution Algorithm (UMDA) [54] is an instance of EDA suited for problems with pseudo-boolean search spaces. UMDA maintains a global probability distribution over the search space, considering the value in every component as a random variable. The core assumption behind UMDA is that those random variables are independent. UMDA is an iterative optimization algorithm. On every iteration, it samples $\lambda$ candidate solutions from the global distribution and uses the first $\mu$ solutions in the ranking according to the objective function to update this probabilistic model. Since usually the algorithm is considered only for pseudo-boolean or continuous domains, we propose our version of UMDA for integer domains in Algorithm 1. In our implementation, we used constants $\mu=5,\lambda=50$.
@@ 4.4 Automated Algorithm Selection
In the previous sections, we introduced several optimization algorithms for SCOP. However, the best choice often depends on subtle features of the considered black-box objective function and can only be determined by empirical testing. To avoid having to run every algorithm on every instance of the OFS problem, we employ Automated Algorithm Selection (AAS) [55], which uses measurable problem properties to recommend or switch between optimizers on the fly. Usually, AAS methods are trained for the specific instances of optimization problems and then applied to previously unseen tasks. In our study, we rely on Nevergrad library and its NGOpt to handle this selection process dynamically. NGOpt is a so-called wizard, which selects an optimization algorithm based on the properties of the considered problem and can switch the algorithm during the minimization process [56]. The properties include the number of dimensions, presence of noise, type of constraints, etc.
@@ 4.5 Numerical Comparison
In this section, we demonstrate the application of the selected algorithms to the OFS problem, see sections 4.14.24.34.4. Let us denote the set of those algorithms as $\mathbb{A}_{1}$. We used $\widehat{\mathrm{F}}_{K}$ as the objective function with fixed constant $K=10^{3}$ for all the algorithms. In each algorithm except BO, we used the computational budget of $2.1\cdot 10^{3}$ evaluations of the objective function $\widehat{\mathrm{F}}_{K}$. For BO we made only $10^{3}$ evaluations of $\widehat{\mathrm{F}}_{K}$, due to the reasons discussed in Section 4.2.1. Every algorithm is run $n=20$ times and every time the set of initial solutions is taken uniformly at random from the search space $\mathscr{L}$, so the initial population is different for each run of every algorithm.
@@
    Quality‐of‐solution.  Quality of the solution considered by algorithm $A$ when computational budget $t$ is spent:
@@
    Best-so-far.  Quality of best-so-far solution found by algorithm $A$ when computational budget $t$ is spent:
@@
We introduce these definitions to be precise about how we measure algorithm performance. Although quality‐of‐solution and best‐so‐far are standard concepts, we did not find their formal definitions in the related literature and so we give them here for clarity and reproducibility.
The results of the numerical experiment are summarized in Figure 3. All the algorithms are given the same total number of the allowed objective function evaluations, except BO, for the reasons discussed in Sec. 4.2.1. In both plots on this figure, the X axis depicts the number of times $t$ function $\widehat{\mathrm{F}}_{K}$ was evaluated. The Y axis in part (a) of Figure 3 represents the approximated mean of quality-of-solution criterion. This plot shows that candidate solutions in Integer Fast Ga, Noisy Portfolio, and BO vary and do not improve on average when the optimization process proceeds. However, NGOpt starts producing solutions that perform better on average when approximately half of the given computational budget is exhausted. This can happen due to the switch of the underlying optimization algorithm that NGOpt uses. Average candidate solutions considered by $(\mu/2,\lambda)$ MIES significantly improve in the first half of the optimization process, but then stagnate until the end.
The Y axis in part (b) of Figure 3 represents the approximated mean of best-so-far criterion and the shaded area represents the approximated standard deviation of best-so-far for each algorithm in $\mathbb{A}_{1}$. The line for every algorithm in this plot constitutes the lower envelope of the corresponding line in the plot (b) of Figure 3. All the algorithms except $(\mu/2,\lambda)$ MIES on average produced some solutions that outperform the baseline solution. It is clear, that algorithm $(\mu+\lambda)$ EA, $(\mu/2+\lambda)$ EA and UMDA outperform all the other considered algorithms in both criteria; quality-of-solution and best-so-far. We call those three algorithms leading algorithms and aim to improve their performance in our further investigation.
@@ 5 Employing New Distances to Navigate in the Domain of OFS
When an optimization algorithm is equipped with domain information it can efficiently disregard areas of the search space with sub-optimal candidate solutions [57]. While there are many possible ways to account for problem-specific information, we use a distance metric as proposed in [57]. The metric represents our knowledge of the filters in the library $\mathbb{L}$ and the search space symmetry to permutations.
5.1 Definition of distance between candidate solutions
In this section, we propose two distance metrics on the space $\mathscr{L}$. We remind the following classical definition of the metric space in B.
@@ 5.1 Definition of distance between candidate solutions
[…]
A larger second moment represents a filter with sharper and or more spectral features. We numerically validated that there are no different filters $x,y\in\mathbb{L}$ such that $d_{1}(x,y)=0$ or $d_{2}(x,y)=0$. The other conditions of the metric are satisfied for $d_{1},d_{2}$ due to the properties of the modulo. Therefore, $d_{1},d_{2}$ are metrics on the space $\mathbb{L}$ (in the mathematical sense).
5.2 Filter multisets
@@
Any two points $\bm{x},\bm{y}\in\mathbb{L}^{M}$ are in relation $\bm{x}\equiv\bm{y}$ if
$\textsc{LAP}(d_{\mathbb{L}},\bm{x},\bm{y})=0.$
This relation induces the quatient space $\mathbb{M}(d_{\mathbb{L}})\coloneqq\mathbb{L}^{M}/\equiv.$
Formally, elements of the space $\mathbb{M}(d_{\mathbb{L}})$ are sets, which are called equivalence classes: $\left\{\bm{x},\bm{y}\in\mathbb{L}^{M}\,\middle|\,\bm{x}\equiv\bm{y}\right\}.$ When we define a metric on the space $\mathbb{M}(d_{\mathbb{L}})$ we show the axioms of metric on points $\bm{x},\bm{y},\bm{z}\in\mathbb{L}^{M}.$ This can be trivially extended to all the elements from the corresponding equivalence classes of $\bm{x},\bm{y},\bm{z}$ using properties of the quotient space. We demonstrate that all the axioms of the metric are upheld in B, therefore $\textsc{LAP}(d_{\mathbb{L}},\cdot,\cdot)$ is metric on the space $\mathbb{M}(d_{\mathbb{L}})$ for any metric $d_{\mathbb{L}}$.
@@ 5.3 Distance-Driven Versions of $(\mu+\lambda)$ EA and $(\mu/2+\lambda)$ EA
Figure 4: Distance-driven optimization. Left: Standard mutation (a) samples points on contours $d_{1},d_{2},d_{3}$ of $F(\bm{x})$; distance-driven mutation (b) “flattens” objective values on closer contours to the reference point $\bm{\widehat{x}}$ in the transformed landscape. Right: To generate a new candidate $\bm{z}$ in the transformed landscape, (c) draw step‐size $s\sim\mathscr{D}(m)$ (see Section 5.3.3), (d) learn scaling parameter $\gamma$ (see Section 5.3.1), (e) rescale the step-size $S={\sqrt{-\ln(1-s)}}/{\gamma},$ (see Section 5.3.1) and (f) solve $\min_{\bm{z}}\bigl|d(\bm{\widehat{x}},\bm{z})-S\bigr|$ to place $\bm{z}$ exactly at distance $S$ from $\bm{\widehat{x}}$ (see Section 5.3.2).
We apply the method proposed by us earlier in [57] to employ the distance metric $d$ in EA. The main steps of this method are visualized in Figure 4. It extends a mutation operator in the optimization algorithm, which aims to find at least one $\bm{x}^{*}\in\mathbb{X}$ such that $F(\bm{x}^{*})=\min\nolimits_{\bm{x}\in\mathbb{X}}\left\{F(\bm{x})\right\},$ where $\mathbb{X}$ is some search space and $F$ is the objective function defined on it. In order to make this extension, the method defines a mapping of the distance between any two candidate solutions to the interval $(0,1)$ and creates a distribution $\mathcal{D}$ over the step-sizes on this interval, visualized in Fig. 4 (c). To do a mutation of the solution $\bm{x}\in\mathbb{X}$, the algorithm samples a step size $s\in(0,1)$ from the distribution $\mathcal{D}$ and maps this step size back to the distance $S\in\mathbb{R}$ in the original domain $\mathbb{X}$, visualized in Fig. 4 (d) and (e). Then it solves the internal optimization problem to generate a solution $\bm{y}\in\mathbb{X}$ which steps away from $\bm{x}$ to the distance $S$:
@@
The landscape of this optimization problem is visualized in Fig. 4 (f). Those steps result in effective transformation of the optimization landscape for every mutation, such that the landscape is flattened on closer contours to the mutated point $\bm{\widehat{x}}$. This is shown in Fig. 4 (a) and (b).
We explain how we applied every step (emphasized with italic font) of this general methodology in later sections: Section 5.3.1, Section 5.3.2, and Section 5.3.3. Now we directly describe the proposed extension of $(\mu/2+\lambda)$ EA summarized in Algorithm 2. The extension of $(\mu+\lambda)$ EA is the same, except that instead of the lines 22, and 2 it chooses $\mathbf{c}$ uniformly at random from $\mathfrak{P}^{(t)}.$
1Select points for exploration for each $i\in\overline{1,R}$: $\mathbf{e}^{(i)}\leftarrow\textsc{u.a.r.}\!\left(\mathbb{L}^{M}\right)$;